OpenStudy (anonymous):

Ok, verifying an identity question 2!

OpenStudy (anonymous):

\[2\sin B \cos ^{3}B + 2\sin ^{3}B \cos B=\sin 2B\] I can't make heads or tails of this!

OpenStudy (turingtest):

The way to do this is not obvious, so I'm going to show you what I will work from to do this problem: http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (turingtest):

so how about try factoring out \(2\sin B\cos B\) from the left and tell me what you get

OpenStudy (anonymous):

That would be something like \[\sin ^{2}B \times \cos ^{2}b\] or would it be sin 2B

OpenStudy (turingtest):

no, I meant\[2\sin B \cos ^{3}B + 2\sin ^{3}B \cos B=2\sin B\cos B(\cos^2B+\sin^2B)\]now how does this simplify?

OpenStudy (anonymous):

The last parenthesis with squared sine and cosine

OpenStudy (anonymous):

is equal to one.

OpenStudy (turingtest):

I was really hoping to let emily realize that on her own juancarlos, but yeah... so why don't you guys go ahead and finish the proof

OpenStudy (anonymous):

You-re right Turing test. sorry.

OpenStudy (turingtest):

it's all good buddy ;) emily is not responding anyway

OpenStudy (anonymous):

ok, it fine guys so\[2\sin B cosB \times(\cos ^{2}B+\sin ^{2}B) or 2\sin B cosB(1) \] but how would I get rid of the cos

OpenStudy (anonymous):

p.s. sorry about the delay, i have a older computer :(

OpenStudy (anonymous):

It should be good if Emily take a look at basic trigonometric identities before attempt to solve any of this "complex" identities.

OpenStudy (anonymous):

lol just realized that was the answer, sorry !

OpenStudy (anonymous):

Thanks both of u, i have helped a lot!

OpenStudy (turingtest):

anytime :D